Given a normal distribution with u = 100 and o = 10, complete parts (a) through (d). Click here to view page 1 of the cumulatz | 0.00 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 -3.0 0.0014 0.0013 0.0013 0.0012 0.0012 0.0011 0.0011 0.0011 0.0010 0.0-1.2|0.1151 0.1131 0.1112 0.1093 0.1075 0.1056 0.1038 0.1020 0.1003 0.0985 -1.110.1357 0.1335 0.13140.1292 0.12710.12510.1230Cumulative standardized normal distribution table (page 2) Z 0.00 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 0.0 0.5000 0.51.7|0.9554 0.9564 0.9573 0.9582 0.9591 0.9599 0.9608 0.9616 0.9625 0.9633 1.870 9641 0 9649 09656 09664 09671 09678 09686 097
Answer #1

a) X (100,10) implies Z=(X-100)/10 N(0,1) P(X>85)=1-(-1.5) = (1.5) = 0.9332 b) P(X<75)=(-2.5) = 0.0062 c) P(X<70 or X>110)=P(

Answers are given in bold. For query, comment.